OpenStudy (anonymous):

How do you find the equation of a line, with m=3, (2, 12) in y=mx+b form?

OpenStudy (anonymous):

a line having slope m and y-intercept b has equation y=mx+b so here y=3x+b as this lline must satisfy the co-ordinates (2,12) so put x=2 and y=12 12=3*2+b 12=6+b b=6 so equation is y=3x+6:)

OpenStudy (anonymous):

@IsabelleLynn96 did u get it????????????

OpenStudy (anonymous):

Yes, and thank you very much. :)